OpenStudy (anonymous):

fan and medal i need help bad A container states that it should be stored at a temperature above twenty degrees Celsius. Convert the temperature to degrees Fahrenheit. Round to the nearest whole number, if necessary. forty-three degrees Fahrenheit fifty-two degrees Fahrenheit sixty-eight degrees Fahrenheit ninety-four degrees Fahrenheit

OpenStudy (fwizbang):

T(F) = 1.8 * T(C) +32

OpenStudy (anonymous):

c/5=(f-32)/9 where c is the temperature value in celcius, f is the temperature value in fahrenheit

OpenStudy (anonymous):

i dont understand :(

OpenStudy (fwizbang):

you know the temperature in Celsius, T(C), so put it the equation and calculate the answer.

OpenStudy (anonymous):

so can u help me with the answer i looked over it over and over again but i still cant get it :'(